An arctic weather balloon is filled with 38.9 L of helium gas inside a prep shed. The temperature inside the shed is 5. oC. The balloon is then taken outside, where the temperature is

−26.°C. Calculate the new volume of the balloon.

You may assume the pressure on the balloon stays constant at exactly 1 atm. Be sure your answer has the correct number of significant digits.
Answer #1

as per Charles' law

V1/T1 = V2/T2 at constant pressure.

V1 = 38.9L

T1 = 5.0oC = 278K

V2 =?

T2 = -26oC= 247K

V2 = V1*T2 /T1

= 38.9L * 247K / 278K = 34.6L

New Volume = 34.6L